HUBBLE SPACE TELESCOPE DAILY REPORT #2868 PERIOD COVERED: 0000Z (UTC) 05/08/01 - 0000Z (UTC) 05/09/01 Daily Status Report as of 129/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 8702 (The Most Elusive Nuclei of LMC Planetary Nebulae) The WF/PC-2 was used to carry out a SNAPSHOT survey of the faintest nuclei of planetary nebulae {PNe} in the LMC. This program will be a crucial follow-on to a Cycle 8 SNAPSHOT of LMC Pne. In thisa case the plan is to observe the faintest or most obscured nuclei to determine their evolutionary state to an accuracy not possible in the Galaxy. These faintest central stars are predicted to be among the most massive nuclei, and owing to the typically advanced age of the surrounding nebulae, are the most sensitive discriminants for validating the post-AGB evolution timescales predicted by theory. The observations completed with no reported problems. 1.2 Completed Ten Sets of STIS/CCD 9248 (Probing the Large Scale Structure: Cosmic Shear Observations) The Space Telescope Imaging Spectrograph (CCD) was used to measure the distortion of light bundles from distant galaxies that probe the statistical properties of the intervening inhomogeneous {dark} matter distribution. Its tidal gravitational field distorts the observable image shapes thereby causing a coherent ellipticity pattern {Cosmic Shear}. As described in yesterday's report and HSTAR 8207, the acquisition for the fourth iteration of this proposal defaulted to fine lock backup on one FGS only, possibly affecting two observations. Then, the loss-of-lock detailed in 2.1 and HSTAR 8210 occurred during the ninth iteration of this proposal, possibly affecting the observations in that iteration. Otherwise, the observations completed nominally. 1.3 Completed Three Sets of WF/PC-2 8828 (Cycle 9 Supplemental Darks pt3/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed with no reported problems. 1.4 Completed Two Sets of STIS/CCD 8864 (CCD Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks for the CCD. The proposal completed nominally. 1.5 Completed Two Sets of STIS/CCD 8845 (Spectroscopic Flats C9) The Space Telescope Imaging Spectrograph (CCD) was used to obtain CCD flats in the spectroscopic mode. The observations completed with no reported problems. 1.6 Completed Two Sets of WF/PC-2 8602 (A Snapshot Survey of the Sites of Recent, Nearby Supernovae) The WF/PC-2 was used to conduct a snapshot survey in V and I of the sites of the nearby SNe, which have precisely known positions, to obtain high-resolution information on their local environment. The proposal completed with no reported problems. 1.7 Completed STIS/CCD 9246 (POMS Test Proposal: Non-scripted Parallel Proposal Continuation IV) The Space Telescope Imaging Spectrograph (CCD) was used to perform a generic target version of the Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The observations completed with no anomalous activity. 1.8 Completed Five Sets of STIS/CCD 8836 (Performance Monitor {Cycle 9}) The Space Telescope Imaging Spectrograph (CCD) was used to measure the baseline performance and commandability of the CCD subsystem. Only primary amplifier D was used. Bias and Flat Field exposures were taken in order to measure read noise, CTE, and gain. Numerous bias frames were taken to permit construction of "superbias" frames in which the effects of read noise have been rendered negligible. Full frame and binned observations were made, with binning factors of 2 x 1, 1 x 2, 2 x 2, 4 x 1, and 4 x 2. Dark images were taken in 2x2 binning mode; 1x1 binning darks are being taken in the nominal CCD Dark Monitor. The exercise completed nominally. 1.9 Completed WF/PC-2 8723 (A Snapshot Survey of Galactic Globular Clusters) The WF/PC-2 was used to observe the crowded centers of globular clusters {GCs} that heretofore have produced intriguing discoveries of new phenomena, most of which are not understood theoretically. The observations completed with no reported problems. 1.10 Completed STIS/CCD 8865 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.11 Completed STIS/CCD 8572 (Identifying Normal Galaxies at 1.3 < z < 2.5) The Space Telescope Imaging Spectrograph (CCD) was used to perform studies of faint, distant galaxies. It is now possible to observe hundreds of galaxies out to z=1 and in the range 3 < z < 4.5, yet the redshift range 1 < z <3 remains largely unexplored. No problems were noted. 1.12 Completed Two Sets of WF/PC-2 8585 (Cosmological Parameters from Type Ia Supernovae at High Redshift) The WF/PC-2 was used to measure the cosmological parameters, Omega, Lambda, and thus the curvature, Omega_k, using Type Ia supernovae {SNe Ia} as calibrated standard candles. As described in yesterday's report and HSTAR 8207, the acquisition for the first iteration of this proposal defaulted to fine lock backup on one FGS only, possibly affecting two observations. Otherwise, there were no other reported problems. 1.13 Completed WF/PC-2 8815 (Cycle 9 Earth Flats) The WF/PC-2 was used to monitor flatfield stability. This proposal obtains sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The proposal completed nominally. 1.14 Completed STIS/CCD/MA2 8569 (A New Survey for Low-Redshift Damped Lyman-Alpha Lines in QSO MgII Systems) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to support studies which have shown that most of the observable neutral gas mass in the Universe resides in QSO damped LyAlpha {DLA} systems. However, at low redshift, DLA can only be found by searching in the UV with HST. By boot-strapping from the MgII statistics, we will be able to further improve the determination of the low- redshift statistical properties of DLA {their incidence and cosmological mass density} and open up new opportunities for studies at low redshift. The observations completed nominally. 1.15 Completed STIS/MA1 8843 (Cycle 9 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to perform the routine monitoring of the MAMA detector dark noise. The proposal completed nominally. 1.16 Completed STIS/CCD 8611 (Observations of Nearby Type Ia Supernovae) The Space Telescope Imaging Spectrograph (CCD) was used to obtain UV spectra of five nearby {0.02 < z < 0.08} SNe Ia in the Hubble Flow. The spectra will be taken at weekly intervals over a range in time starting slightly before maximum light and extending to +30 days. These observations will accomplish the following three goals: {1} calibration of the rest frame UV light curves of SNe Ia and an assessment of their potential use as distance indicators through UV light curve shape analyses. {2} improvement in our understanding of the physics of SNe Ia, metallicity/evolutionary effects and correlations between peak brightness and UV spectral features. {3} calibration of the SNe Ia previously observed by HST at high-redshift. This data is crucial for proper cross-filter k-corrections and calibration of the supernova photometry. The observations completed with no reported anomalous activity. 1.17 Completed Two Sets of WF/PC-2 8604 (Stellar Populations Across the Small Magellanic Cloud (SMC) : History and Structure) The WF/PC-2 was used to investigate the structure and evolutionary history of the SMC by obtaining a series of 50 three-color snapshots of selected SMC regions. With this polling of SMC field star properties, we will produce color-magnitude diagrams in the U, V and I bands which will reach V~23.5 in regions too crowded to be observed accurately from the ground. The observations completed nominally. 1.18 Completed WF/PC-2 8682 (A Snapshot Study of 0bservational Cosmology) The WF/PC-2 was used to examine the observational constraints on the cosmic star formation history that is currently among the most active fields in observational cosmology. The most widely used tracer of the co-moving volume-averaged star formation rate {SFR} is the UV luminosity density, which early results found to peak at z~1- 2. The apparent identification of the primary epoch of metal production and star formation in the Universe led to intense theoretical and observational interest. Nevertheless, and remarkably for such a fundamental observation, little is known about the history of star formation in the Universe beyond its global average. The loss-of-lock detailed in 2.1 and HSTAR 8210 occurred during this proposal, possibly affecting the observations. Otherwise, there were no other reported anomalies. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 12 Successful: 12 As described in HSTAR 8210, there was a loss-of-lock on FGS-3 at 128/223631Z after a successful acquisition at 128/222954Z. The proposals presented in 1.2 and 1.18 may have been affected. Scheduled Re-acquisitions: 4 Successful: 4 2.2 FHST Updates: Scheduled: 23 Successful: 23 2.3 Operations Notes: Using ROP SR-1A, the SSR EDAC error counter was cleared three times. Per an operations request, gyro #5 was successfully turned on at 128/1210Z. The initial motor signature current and rate output were as expected. The motor current displayed an itial spike and then settled to a nominal value of 140 mAmps. At 128/1515Z, the motor current began an upward climb, necessitating several modifications of the limits for that parameter. Then, at 129/041654Z, the motor current rose to 250 mAmps, and over a period of a minute continued to rise to the stall current of 320 mAmps. Shortly after a forward service was acquired, flight operations personnel powered off gyro #5 at 129/042056Z per an operations request. The spinning restart was not executed since the gyro has already been above the 320 mAmps level for three minutes by the time that the forward service occurred. The CCS status buffer was dumped and cleared at 128/1538Z per ROP DF-18A. Using the same ROP, the CCS status buffer limits were adjusted at 1282315Z. The NSSC-1 status buffer was dumped and reset at 128/2323Z per ROP NS-3. A TTR was written for a required re-transmit during a NSSC-1 load uplink at 129/0330Z. The NSSC-1 engineering report #0 was reset at 129/0341Z as directed by ROP NS-5. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW